OpenStudy (anonymous):

graph the system of inequalities x + 3y<= 9, 3x + y <= 9, x >= 0, y >= 0

OpenStudy (anonymous):

dumbcow can you help me with this one please

OpenStudy (anonymous):

treat them like simultaneous equations multriply second equation by 3 to get 9x + 3y <= 27 takes this away from forst equation, to get: (x-9x) + (3y-3y) = 9-27, so -8x <= -18, so x <= 9/4 and so on...